For the function f(x) = x (squared) - 6x +2
What is the y-intercept? (I got -7)
What is the vertex? (I got 3, -7)
I'm doing homework and feel like my answers are wrong. Please find the answers and let me know so I can check my work. Thank you =)

Answer by Alan3354(69443)

For the function f(x) = x^2 - 6x +2
What is the y-intercept? (I got -7)
The y-intercept is at x = 0
f(0) = 0 - 0 + 2
= 2
=================
What is the vertex? (I got 3,-7)
--------------
The vertex is on the Axis of Symmetry.
The AOS is x = -b/2a
x = 6/4 = 3/2
f(3/2) = (9/4) - 9 + 2 = -19/4
Vertex at (3/2,-19/4)